# Variational Search Distributions

### v2

@@ 2.4 Theoretical Analysis
In this section we summarize the main theoretical results concerning VSD and its estimates. We show that VSD sampling distributions converge to a target distribution that characterizes the level set given by $\tau$, satisfying (D1) Guaranteed in two general settings. We first derive results assuming ${f\!\centerdot}$ is drawn from a Gaussian process, i.e., ${f\!\centerdot}\sim\mathcal{GP}(0,k)$, with a positive-semidefinite covariance (or kernel) function $k:\mathcal{X}\times\mathcal{X}\to$ (Appendix D) and then using GP-PI as the CPE for VSD. These results are then extended to probabilistic classifiers based on wide neural networks (Appendix E) by means of the neural tangent kernel (NTK) for the given architecture (Jacot et al., 2018). For simplicity, we set $B=1$ and $N=t$, though sampling with $B>1$ should improve rates by a multiplicative factor. {restatable}theoremmainthm Let assumptions D.1 to D.5 hold. Then VSD equipped with GP-PI approaches the level-set distribution at the following rate:
@@
This result is based on showing that the GP posterior variance vanishes at an optimal rate of ${\mathcal{{O}}}(t^{-1})$ in our setting (D.5). We also analyze the rate at which VSD finds feasible designs, or “hits”, compared to an oracle with full knowledge of ${f\!\centerdot}$. After $T$ rounds, the number of hits found by VSD is $H_{T}=\sum_{t=1}^{T}\mathbbm{1}[{y_{t}>\tau_{t-1}}]$, where $y_{t}$ is generated from Equation 1 and $\mathbf{x}_{t}\sim p({\mathbf{x}}|{y>\tau_{t-1},\mathcal{D}_{t-1}})$. The number of hits, $H_{T}^{*}$, from an agent that fully knows ${f\!\centerdot}$ is the same but for generating conditioned on ${f\!\centerdot}$ with $\mathbf{x}_{t}\sim p({\mathbf{x}}|{y>\tau_{t-1},{f\!\centerdot}})$. Using this definition and Sec. 2.4 we show the following. {restatable}corollaryhitscor Under the settings in Sec. 2.4, we also have that:
@@
$\mathbb{E}[H_{T}]$ is related to the empirical recall measure in Equation 16 up to the normalization constant, but it does not account for repeated hits, which are treated as false discoveries (false positives) under recall. Lastly, for NN-based CPEs, we obtain convergence rates dependent on the spectrum of the NTK (E.2), which we instantiate for ReLU networks below. For the full results and proofs, please see Appendix D for the GP-based analysis and Appendix E for the NTK results. {restatable}corollaryrelukl Let $\pi_{\theta}$ be modeled via a fully connected ReLU network. Then, under the assumptions in E.2, VSD achieves:
$\mathbb{D}\!\left[{p({\mathbf{x}}|{y>\tau_{t},\mathcal{D}_{t}})}\|{p({\mathbf{x}}|{y>\tau,{f\!\centerdot}})}\right]\in\widetilde{{\mathcal{{O}}}}_{\mathbb{P}}\left(t^{-\frac{1}{2(M+1)}}\right),$ (12)
which asymptotically vanishes for all finite sequence lengths $M$.
@@ 3 Related Work
VSD can be viewed as one of many methods that makes use of the bound (Staines & Barber, 2013),
@@
The maximum is always greater than or equal to the expected value of a random variable. This bound is useful for black-box optimization (BBO)of ${f\!\centerdot}$, and becomes tight if $q({\mathbf{x}}|{\phi})\to\delta(\mathbf{x}^{*})$, see Appendix F for more detail and VSD’s relation. Other well known methods that make use of this bound are evolution strategies (ES) and natural evolution strategies (NES) (Wierstra et al., 2014), variational optimization (VO) (Staines & Barber, 2013; Bird et al., 2018), estimation of distribution algorithms (EDA) (Larrañaga & Lozano, 2001), and Bayesian optimization with probabilistic reparameterisation (BOPR) (Daulton et al., 2022). For learning the parameters of the variational distribution, $\phi$, they variously make use of maximum likelihood estimation or the score function gradient estimator (REINFORCE) (Williams, 1992). Algorithms that modify Equation 13 to stop the collapse of $q({\mathbf{x}}|{\phi})$ to a point mass for batch design include design by adaptive sampling (DbAS) (Brookes & Listgarten, 2018) and conditioning by adaptive sampling (CbAS) (Brookes et al., 2019). They use fixed samples $\mathbf{x}^{(s)}$ from $q({\mathbf{x}}|{\phi^{*}_{t-1}})$ for approximating the expectation, and then optimize $\phi$ using a weighted maximum-likelihood or variational style procedure. We can take a unifying view of many of these algorithms by recognizing the general gradient estimator,
$\mathbb{E}_{q({\mathbf{x}}|{\phi^{\prime}})}\!\left[{w(\mathbf{x})\nabla_{\phi}\log q({\mathbf{x}}|{\phi})}\right],$ (14)
where we give each component in Table 1. BORE has been adapted to discrete $\mathcal{X}$ by using the score function gradient estimator and CbAS and DbAS have been adapted to use a CPE– their original derivations use the equivalent of a PI acquisition function.
Method $w(\mathbf{x})$ $\phi^{\prime}$ Fixed $\mathbf{x}^{(s)}\sim q({\mathbf{x}}|{\phi^{\prime}})$?
VSD $\log\pi_{\theta^{*}}({\mathbf{x}})+\log p({\mathbf{x}}|{\mathcal{D}_{0}})-\log q({\mathbf{x}}|{\phi})$ $\phi$ No
CbAS $\pi_{\theta^{*}}({\mathbf{x}})p({\mathbf{x}}|{\mathcal{D}_{0}})/q({\mathbf{x}}|{\phi^{*}_{t-1}})$ $\phi^{*}_{t-1}$ Yes
DbAS $\pi_{\theta^{*}}({\mathbf{x}})$ $\phi^{*}_{t-1}$ Yes
BORE $\pi_{\theta^{*}}({\mathbf{x}})$ $\phi$ No
BOPR $\alpha\!\left({\mathbf{x},\mathcal{D}_{N}}\right)$ $\phi$ No
@@
A number of finite horizon methods have been applied to biological sequence BBO tasks, such as Amortized BO (Swersky et al., 2020), GFlowNets (Jain et al., 2022), and the reinforcement learning based DynaPPO (Angermueller et al., 2019). Heuristic stochastic search methods such as AdaLead (Sinai et al., 2020) and proximal exploration (PEX) (Ren et al., 2022) have also demonstrated strong empirical performance on these tasks. We compare the properties of the most relevant methods to our problem in Table 2.
Other methods that explicitly attempt to estimate the super level-set distribution, $p({\mathbf{x}}|{y>\tau})$, include CbAS, which optimizes the forward KL divergence, $\mathbb{D}_{\mathrm{KL}}\!\left[{p({\mathbf{x}}|{y>\tau})}\|{q({\mathbf{x}}|{\phi})}\right]$, in a sequential fashion using importance weighted cross entropy estimation (Rubinstein, 1999). Batch-BORE (Oliveira et al., 2022) also optimizes the reverse KL divergence and uses CPE, but with Stein variational inference (Liu & Wang, 2016) for continuous and diverse batch candidates $\{\mathbf{x}_{tb}\in^{M}\}^{B}_{b=1}$ (the authors optimize on the simplex, and then choose the maximum coordinate for their discrete experiments). There is a rich literature on the related task of active learning and BO for level set estimation (LSE) (Bryan et al., 2005; Gotovos et al., 2013; Bogunovic et al., 2016; Zhang et al., 2023). However we focus on the task of learning a generative model over a practically innumerable discrete space.
Method     Rare $\mathbf{x}\in\mathcal{S}$ (R1) Rare     Sequential (R2) Sequential     Discrete $\mathcal{X}$ (R3) Discrete     Batch $\{\mathbf{x}_{bt}\}^{B}_{b=1}$ (R4) Batch     Guaranteed (D1) Guaranteed     Gradient descent (D2) Gradient     Generative $q({\mathbf{x}}|{\phi})$ (D3) Generative     Scalable (D4) Scalable     General acq./reward fn.     Amortization
@@
Table 2: Feature table of competing methods: ✓ has feature, ✗ does not have feature, – partially has feature, or requires simple modification. We follow Swersky et al. (2020) in their definition of amortization referring to the ability to use $q({\mathbf{x}}|{\phi^{*}_{t-1}})$ for warm-starting the optimization of $\phi_{t}$.
@@ 4 Experiments
Firstly we test our method, VSD, on its ability to generate complex, structured candidates, $\mathbf{x}$, in a single round by training it to generate a subset of handwritten digits from flattened MNIST images (LeCun et al., 1998) in Sec. 4.1. We then compare VSD on two sequence design tasks against existing baseline methods. The first of these tasks (Sec. 4.2) is to discover as many unique, fit sequences as possible using the datasets DHFR (Papkou et al., 2023), TrpB (Johnston et al., 2024) and TFBIND8 (Barrera et al., 2016). These datasets contain near complete evaluations of $\mathcal{X}$, and to our knowledge DHFR and TrpB are novel in the machine learning literature. The second (Sec. 4.3) is a more traditional black-box optimization task of finding the maximum of an unknown function; using datasets AAV (Bryant et al., 2021) and GFP (Sarkisyan et al., 2016). A complete combinatorial assessment is infeasible for these latter datasets, and so we use the convolution neural network oracle presented in Kirjner et al. (2024) as in-silico ground truth. The corresponding datasets involve $|\mathcal{V}|\in\{4,20\}$, $4\leq M\leq 237$ and $65,000<|\mathcal{X}|<20^{237}$. We discuss the settings and properties of these datasets in greater detail in Appendix B.
For the biological sequence experiments we run a predetermined number of experimental rounds, $T=10$, we set the batch size to $B=128$, and use five different seeds for random initialization. For the first set of real-sequence experiments in Sec. 4.2, we use a fixed threshold, $\tau$, with the aim of finding $\mathbf{x}\in\mathcal{S}$ using both probabilistic and non probabilistic models. For the next set of experiments in Sec. 4.3, we set $\tau_{t}$ adaptively each round for testing VSD’s ability to find the maximizer, $\mathbf{x}^{*}$, of the black-box function, ${f\!\centerdot}$. We compare against DbAS (Brookes & Listgarten, 2018), CbAS (Brookes et al., 2019), AdaLead (Sinai et al., 2020), PEX (Ren et al., 2022), BORE (Tiao et al., 2021) adapted to use the score function gradient estimator, and a naïve baseline that uses random samples from the prior, $p({\mathbf{x}}|{\mathcal{D}_{0}})$. To reduce confounding, all methods share the same surrogate model and other components where possible (acquisition functions, priors and variational distributions). We optimize VSD, CbAS, DbAS and BORE for a minimum of 5000 iterations each round. When we use a CPE, AdaLead’s $\kappa$ parameter is set to 0.5 since the CPE already incorporates the appropriate threshold.
@@ 4.1 Conditional Generation of Handwritten Digits
Our motivating application for VSD is to model the space of fit DNA and protein sequences, which are string-representations of complex 3-dimensional structures. In this experiment we aim to demonstrate, by analogy, that VSD can operate with the types of generative models required for representing these complex structures. For this task, we have chosen to ‘unroll’ (reverse the order of every odd row, and flatten) down-scaled ($14\times 14$ pixel, 8-bit) MNIST (LeCun et al., 1998) images into sequences, $\mathbf{x}$, where $M$ = 196 and $|\mathcal{V}|=8$. We then train long short-term memory (LSTM) recurrent neural network (RNN) and decoder-only causal transformer generative models on the entire MNIST training set by maximum likelihood (ML). These generative distributions are used as our prior models, $p({\mathbf{x}}|{\mathcal{D}_{0}})$, for VSD and we detail their form in Sec. B.1.
@@
The task is then to use VSD in one round to estimate the posterior $p({\mathbf{x}}|{y\in\{3,5\}})$ using a CPE trained on labels $z_{n}=\mathbbm{1}[{y_{n}\in\{3,5\}}]$. We use a convolutional architecture for the CPE given in Sec. B.2, and it achieves a test balanced accuracy score of $\sim 99\%$. We parameterize the variational distributions, $q({\mathbf{x}}|{\phi})$, in the same way as the priors, and initialize these distribution parameters from the prior distribution parameters. During training with ELBO the prior distribution parameters are locked, and we run training for 5000 iterations. This is exactly lines 8 and 9 in Algorithm 1. Samples are visualized from the resulting variational distributions with the corresponding priors in Figure 2. We can see that the prior LSTM and transformer are able to generate convincing digits once the sampled sequences are ‘re-rolled’, and that VSD is able to effectively refine these distributions, even though it does not have access to any data directly – only scores from the CPE. Both the LSTM and transformer yield qualitatively similar results, and have similar mean scores from the CPE.

### v4

@@ 2.4 Theoretical Analysis
We show that VSD sampling distributions converge to a target distribution that characterizes the level set given by $\tau$, satisfying (D1) Guaranteed in two general settings. We first derive results assuming ${f\!\centerdot}$ is drawn from a Gaussian process, i.e., ${f\!\centerdot}\sim\mathcal{GP}(0,k)$, with a positive-semidefinite covariance (or kernel) function $k:\mathcal{X}\times\mathcal{X}\to$ (Appendix E), using GP-PI as the CPE for VSD. These results are then extended to probabilistic classifiers based on wide neural networks (NNs) (Appendix F) by means of the neural tangent kernel (NTK) for the given architecture (Jacot et al., 2018). For the analysis, we set $B=1$ and $N=t$, though having $B>1$ should improve the rates by a multiplicative factor. {restatable}theoremmainthm Under mild assumptions (E.1 to E.5), the variational distribution of VSD equipped with GP-PI converges to the level-set distribution in probability at the following rate:
@@
This result is based on showing that the GP posterior variance vanishes at an optimal rate of ${\mathcal{{O}}}(t^{-1})$ in our setting (E.5). We also analyze the rate at which VSD finds feasible designs, or “hits”, compared to an oracle with full knowledge of ${f\!\centerdot}$. After $T$ rounds, the number of hits found by VSD is $H_{T}=\sum_{t=1}^{T}\mathds{1}[{y_{t}>\tau_{t-1}}]$, where $y_{t}$ follows Equation 1 and $\mathbf{x}_{t}\sim p({\mathbf{x}}|{y>\tau_{t-1},\mathcal{D}_{t-1}})$. The number of hits, $H_{T}^{*}$, from an agent that fully knows ${f\!\centerdot}$ is the same but for generating conditioned on ${f\!\centerdot}$ with $\mathbf{x}_{t}\sim p({\mathbf{x}}|{y>\tau_{t-1},{f\!\centerdot}})$. Using this definition and Sec. 2.4, we prove the following. {restatable}corollaryhitscor Under the settings in Sec. 2.4, we also have that:
@@
$\mathbb{E}[H_{T}]$ is related to the empirical recall measure (16) up to the normalization constant, but it does not account for repeated hits, which are treated as false discoveries (false positives) under recall. Lastly, for NN-based CPEs, we obtain convergence rates dependent on the spectrum of the NTK (F.2), which we instantiate for infinitely wide ReLU networks below. For the full results and proofs, please see Appendix E for the GP-based analysis and Appendix F for the NTK results. {restatable}corollaryrelukl Let $\pi_{\theta}$ be modeled via a fully connected ReLU network. Then, under assumptions on identifiability and sampling (F.1 to F.6), in the infinite-width limit, VSD with CPE-PI achieves:
$\mathbb{D}\!\left[{p({\mathbf{x}}|{y>\tau_{t},\mathcal{D}_{t}})}\|{p({\mathbf{x}}|{y>\tau,{f\!\centerdot}})}\right]\in\widetilde{{\mathcal{{O}}}}_{\mathbb{P}}\left(t^{-\frac{1}{2(M+1)}}\right).$ (12)
This result finally indicates that, when equipped with flexible NN-based CPEs, VSD is also capable of recovering the target distribution for arbitrary sequence lengths in combinatorial problems.
@@ 3 Related Work
We will consider related work firstly in terms of methods that have similar components to VSD, then secondly in terms of related problems to our specification of active generation. VSD can be viewed as one of many methods that makes use of the variational bound (Staines & Barber, 2013),
@@
The maximum is always greater than or equal to the expected value of a random variable. This bound is useful for black-box optimization (BBO)of ${f\!\centerdot}$, and becomes tight if $q({\mathbf{x}}|{\phi})\to\delta(\mathbf{x}^{*})$. See Appendix G for more detail and VSD’s relation to BO. Other well known methods that make use of this bound are natural evolution strategies (NES) (Wierstra et al., 2014), variational optimization (VO) (Staines & Barber, 2013; Bird et al., 2018), estimation of distribution algorithms (EDA) (Larrañaga & Lozano, 2001; Brookes et al., 2020), and Bayesian optimization with probabilistic reparametrization (BOPR) (Daulton et al., 2022). For learning the parameters of the variational distribution, $\phi$, they variously make use of maximum likelihood estimation or the score function gradient estimator (REINFORCE) (Williams, 1992). Algorithms that explicitly modify Equation 13 to stop the collapse of $q({\mathbf{x}}|{\phi})$ to a point mass for batch design include design by adaptive sampling (DbAS) (Brookes & Listgarten, 2018) and conditioning by adaptive sampling (CbAS) (Brookes et al., 2019). They use fixed samples $\mathbf{x}^{(s)}$ from $q({\mathbf{x}}|{\phi^{*}_{t-1}})$ for approximating the expectation, and then optimize $\phi$ using a weighted maximum-likelihood or variational style procedure. Though DbAS and CbAS were formulated for offline (non-sequential) tasks, they have often been used in a sequential setting (Ren et al., 2022). We can take a unifying view of algorithms that use a surrogate model for ${f\!\centerdot}$ by recognizing the general gradient estimator,
$\mathbb{E}_{q({\mathbf{x}}|{\phi^{\prime}})}\!\left[{w(\mathbf{x})\nabla_{\phi}\log q({\mathbf{x}}|{\phi})}\right].$ (14)
where we give each component in Table 1. For our experiments BORE has been adapted to discrete $\mathcal{X}$ by using the score function gradient estimator, which we denote by BORE , while CbAS and DbAS have been adapted to use a CPE– their original derivations use a PI acquisition function.
Method $w(\mathbf{x})$ $\phi^{\prime}$ Fixed $\mathbf{x}^{(s)}\sim q({\mathbf{x}}|{\phi^{\prime}})$ per round?
VSD $\log\pi_{\theta^{*}}({\mathbf{x}})+\log p({\mathbf{x}}|{\mathcal{D}_{0}})-\log q({\mathbf{x}}|{\phi})$ $\phi$ No (REINFORCE)
CbAS $\pi_{\theta^{*}}({\mathbf{x}})p({\mathbf{x}}|{\mathcal{D}_{0}})/q({\mathbf{x}}|{\phi^{*}_{t-1}})$ $\phi^{*}_{t-1}$ Yes (importance Monte Carlo)
DbAS $\pi_{\theta^{*}}({\mathbf{x}})$ $\phi^{*}_{t-1}$ Yes (Monte Carlo)
BORE $\pi_{\theta^{*}}({\mathbf{x}})$ $\phi$ No (REINFORCE)
BOPR $\alpha\!\left({\mathbf{x},\mathcal{D}_{N}}\right)$ $\phi$ No (REINFORCE)
@@
A number of finite horizon methods have been applied to biological sequence BBO tasks, such as Amortized BO (Swersky et al., 2020), GFlowNets (Jain et al., 2022), and the reinforcement learning based DynaPPO (Angermueller et al., 2019). LSO-like methods (Gómez-Bombarelli et al., 2018; Tripp et al., 2020; Stanton et al., 2022; Gruver et al., 2023) tackle optimization of sequences by encoding them into a continuous latent space within which candidate optimization or generation takes place. Selected candidates are decoded back into sequences before black box evaluation; see González-Duque et al. (2024) for a comprehensive survey. VSD does not require a latent space nor an encoder, and as such can be seen as an amortized variant of probabilistic reparameterisation methods (Daulton et al., 2022) or continuous relaxations (Michael et al., 2024). Heuristic stochastic search methods such as AdaLead (Sinai et al., 2020) and proximal exploration (PEX) (Ren et al., 2022) have also demonstrated strong empirical performance on these tasks. We compare the properties of the most relevant methods to our problem in Table 2.
In contrast to just finding the maximum using BBO, active generation considers another problem – generating samples from a rare set of feasible solutions. Generation methods that estimate the super level-set distribution, $p({\mathbf{x}}|{y>\tau})$, include CbAS, which optimizes the forward KL divergence, $\mathbb{D}_{\mathrm{KL}}\!\left[{p({\mathbf{x}}|{y>\tau})}\|{q({\mathbf{x}}|{\phi})}\right]$ using importance weighted cross entropy estimation (Rubinstein, 1999). Batch-BORE (Oliveira et al., 2022) also optimizes the reverse KL divergence and uses CPE, but with Stein variational inference (Liu & Wang, 2016) for diverse batch candidates (with a continuous relaxation for discrete variables). There is a rich literature on the related task of active learning and BO for level set estimation (LSE) (Bryan et al., 2005; Gotovos et al., 2013; Bogunovic et al., 2016; Zhang et al., 2023a). However, we focus on learning a generative model of a discrete space.
For active generation VSD, CbAS and DbAS all use an acquisition function defined in the original domain, $\mathcal{X}$, to weight gradients (see Equation 14) for learning a conditional generative model, from which $\mathbf{x}_{bt}$ are sampled. An alternative is to use guided generation, that is to train an unconditional generative model, and then have a discriminative model guide (condition) the samples from the unconditional model at test time. This plug-and-play of a discriminative model has shown promise for controlled image and text generation of pre-trained models (Nguyen et al., 2017; Dathathri et al., 2020; Li et al., 2022; Zhang et al., 2023b). LaMBO (Stanton et al., 2022) and LaMBO-2 (Gruver et al., 2023) take a guided generation approach to solve the active generation problem. LaMBO uses an (unconditional) masked language model auto-encoder, and then optimizes sampling from its latent space using an acquisition function as a guide. LaMBO-2 takes a similar approach, but uses a diffusion process as the unconditional model, and modifies a Langevin sampling de-noising process with an acquisition function guide.
Method     Rare $\mathbf{x}\in\mathcal{S}$ (R1) Rare     Sequential (R2) Sequential     Discrete $\mathcal{X}$ (R3) Discrete     Batch $\{\mathbf{x}_{bt}\}^{B}_{b=1}$ (R4) Batch     Generative $q({\mathbf{x}}|{\phi})$ (R5) Generative     Guaranteed (D1) Guaranteed     Gradient descent (D2) Gradient     Scalable (D3) Scalable     General acq./reward fn.     Amortization
@@
Table 2: Feature table of competing methods: ✓ has feature, ✗ does not have feature, – partially has feature, or requires only simple modification. We follow Swersky et al. (2020) in their definition of amortization referring to the ability to use $q({\mathbf{x}}|{\phi^{*}_{t-1}})$ for warm-starting the optimization of $\phi_{t}$.
@@ 4 Experiments
Firstly we test our method, VSD, on its ability to generate complex, structured designs, $\mathbf{x}$, in a single round by training it to generate a subset of handwritten digits from flattened MNIST images (LeCun et al., 1998) in Sec. 4.1. We then compare VSD on two sequence design tasks against existing baseline methods. The first of these tasks (Sec. 4.2) is to generate as many unique, fit sequences as possible using the datasets DHFR (Papkou et al., 2023), TrpB (Johnston et al., 2024) and TFBIND8 (Barrera et al., 2016). These datasets contain near complete evaluations of $\mathcal{X}$, and to our knowledge DHFR and TrpB are novel in the machine learning literature. The second (Sec. 4.3) is a more traditional black-box optimization task of finding the maximum of an unknown function; using datasets AAV (Bryant et al., 2021), GFP (Sarkisyan et al., 2016) and the biologically inspired Ehrlich functions (Stanton et al., 2024). The corresponding datasets involve $|\mathcal{V}|\in\{4,20\}$, $4\leq M\leq 237$ and $65,000<|\mathcal{X}|<20^{237}$. We discuss the settings and properties of these datasets in greater detail in Appendix C. For the biological sequence experiments we run a predetermined number of experimental rounds, $T=10$ or $32$ for the Ehrlich functions. We set the batch size to $B=128$, and use five different seeds for random initialization. We compare against DbAS (Brookes & Listgarten, 2018), CbAS (Brookes et al., 2019), AdaLead (Sinai et al., 2020), and PEX (Ren et al., 2022) – all of which we have adapted to use a CPE, BORE (Tiao et al., 2021) – which we have adapted to use the score function gradient estimator, and a naïve baseline that uses random samples from the prior, $p({\mathbf{x}}|{\mathcal{D}_{0}})$. To reduce confounding, all methods share the same surrogate model, acquisition functions, priors and variational distributions where possible. We compare against LaMBO-2 (Gruver et al., 2023) on the Ehrlich functions, it uses its own surrogate and generative models.
@@ 4.1 Conditional Generation of Handwritten Digits
Our motivating application for VSD is to model the space of fit DNA and protein sequences, which are string-representations of complex 3-dimensional structures. In this experiment we aim to demonstrate, by analogy, that VSD can generate sequences that represent 2-dimensional structures. For this task, we have chosen to ‘unroll’ (reverse the order of every odd row, and flatten) down-scaled ($14\times 14$ pixel, 8-bit) MNIST (LeCun et al., 1998) images into sequences, $\mathbf{x}$, where $M$ = 196 and $|\mathcal{V}|=8$. We then train long short-term memory (LSTM) recurrent neural network (RNN) and decoder-only causal transformer generative models on the entire MNIST training set by maximum likelihood (ML). These generative distributions are used as the prior models, $p({\mathbf{x}}|{\mathcal{D}_{0}})$, for VSD and we detail their form in Appendix C.3.
@@
The task is then to use VSD in one round to estimate the posterior $p({\mathbf{x}}|{y\in\{3,5\}})$ using a CPE trained on labels $z_{n}=\mathds{1}[{y_{n}\in\{3,5\}}]$. We use a convolutional architecture for the CPE given in Appendix C.4, and it achieves a test balanced accuracy score of $\sim 99\%$. We parameterize the variational distributions, $q({\mathbf{x}}|{\phi})$, in the same way as the priors, and initialize these distribution parameters from the prior distribution parameters. During training with ELBO the prior distribution parameters are locked, and we run training for 5000 iterations. This is exactly lines 8 and 9 in Algorithm 1. Samples are visualized from the resulting variational distributions with the corresponding priors in Figure 2. We see that the prior LSTM and transformer are able to generate convincing digits once the sampled sequences are ‘re-rolled’, and that VSD is able to effectively refine these distributions, even though it does not have access to any data directly – only scores from the CPE. Both the LSTM and transformer yield qualitatively similar results, and have similar mean scores from the CPE.